# CzechLynx: A Dataset for Individual Identification and Pose Estimation of the Eurasian Lynx

### v1

@@ 2.5 Synthetic Data Generation
Synthetic data has proven valuable in improving performance in many scenarios, especially when real-world data is scarce, difficult to collect, or lacks sufficient diversity [32, 8, 33]. In the context of individual animal identification or pose estimation, however, generating synthetic data that accurately captures the complexity of real-world conditions remains a significant challenge. Despite its usefulness, existing methods for generating synthetic animal pose data often struggle with realism in appearance, motion, and environmental context, which can lead to reduced accuracy in downstream applications [34, 35, 36]. To complement the real data and broaden the variability in both poses and textures, we created a synthetic dataset using rendered 2D images of a 3D lynx model. Our generation pipeline, built with the Unity game engine, enables the creation of highly realistic synthetic samples that reflect multiple individuals of the Lynx lynx species across a wide range of conditions. By leveraging detailed environmental modeling, including vegetation, terrain, and lighting, along with lifelike animations and diverse camera setups, we ensure the production of high-fidelity synthetic images that resemble real-world scenarios.
1:Blank 3D mesh of animal, set of textures (each with probability of being selected $p_{t}$), set of animations (each with probability of being selected $p_{a}$)
@@
9:             With probability $p_{t}$ select a texture for the 3D mesh
10:             With probability $p_{a}$ trigger auxiliary animations
11:             Capture and store a 2D snapshot every $N$-th frame
Pipeline: We propose a pipeline to generate synthetic animal images with rich metadata, including precise pose, segmentation masks, and individual annotations. The process begins with the construction of a 3D environment in Unity, where we simulate various behavioral scenarios. This involves modeling the animal in 3D with high-resolution textures and animations, as well as adjusting environmental elements to introduce naturalistic variability. We include an animal moving through a predefined environment using a walking animation that follows multiple fixed points. An auxiliary sitting animation is randomly triggered approximately once per path between two points to include some pose variability. A single camera captures 2D snapshots of the scene, recording metadata such as pose keypoints. The camera placement ensures that all animal keypoints remain visible and unobstructed throughout the simulation. By integrating these components, we create a dynamic simulation that closely mimics real-world animal behavior. To generate multiple identities, each frame uses a randomly selected lynx texture, with each texture corresponding to a unique individual. The complete procedure is outlined in Algorithm 11.
Data generation pipeline: Our data generation process is designed to closely replicate real-world camera trap photography. We begin by manually selecting four points in each scene, some of which may be outside the camera’s field of view. The animal model then moves randomly between these points. As it walks, we randomly select $n$stop" points along the path. At each "stop", we modify both the scene and the model’s pose. Scene adjustments include selecting different environment variations (e.g., tree types, grass, snow) and making slight changes to the camera viewpoint (e.g., rotation). The model’s pose is altered by rotating it and transitioning between available animations. If a texture variation is provided, it is also randomly applied. We save two data samples: one before any adjustments and another after modifications. If the majority of the model is outside the camera’s view, the sample is discarded.
@@
Animal model: We use a textured model of Lynx lynx as our animal model. With proper setup, other animal models can also be used. Unfortunately, the model initially included only a walking animation. To increase pose diversity, we manually created an additional sitting animation. The model’s skeletal structure includes 20 keypoints and follows the Animal Pose dataset. This allows for training on both real and synthetic datasets using the same model and enables direct performance comparisons. Additionally, we capture the model’s bounding box, inferred from its mesh.
Textures: For the 3D animal meshes used in the simulation, we produce an alternative set of textures. Rather than manually constructing each texture, we use diffusion-based models [37, 38, 39, 40], which have proven highly effective for various computer vision tasks, including texture generation for 3D meshes and enables the creation of textures directly based on text descriptions. In our pipeline, we employ the Paint-It model [40], a text-to-texture model that generates textures for blank 3D meshes based on textual prompts. To generate realistic and varied textures for our models, we use prompts such as lynx and bobcat, which are often used interchangeably in the context of image labels. Using these prompts allows the model to produce different textures that reflect the unique features of these animals. However, as shown in the first row of Figure 3, these basic prompts did not always result in diverse textures. To increase the variability, we introduced an additional 24 prompts – 12 for lynx and 12 for bobcat. These prompts provided detailed descriptions of common feline fur patterns, incorporating terms such as spotted, tabby, and marbled. As shown in the second row of Figure 3, the inclusion of these detailed prompts significantly increased the variability of the generated textures without causing significant qualitative distortion. Overall, we produced 300 unique textures.
# Samples Full text prompt
15+15 A lynx / bobcat.
$15+15$ A lynx / bobcat in natural habitat.
$10+10$ A Spotted lynx / bobcat with small, evenly spaced black spots scattered across its beige fur, showcasing a classic spotted pattern.
$10+10$ A Striped lynx / bobcat featuring thin, dark stripes running along its back and sides, adding a sleek and striking appearance to its coat.
$10+10$ A Marbled lynx / bobcat with a marbled pattern, characterized by swirling patches of lighter and darker fur, creating a mesmerizing and intricate design.
$10+10$ A Tabby lynx / bobcat displaying tabby markings, with bold, dark stripes running down its sides and a distinctive shape on its forehead, giving it a distinctive and recognizable look.
$10+10$ A Rosette lynx / bobcat with rosette-shaped markings, consisting of dark outlines surrounding clusters of lighter fur, resembling the spots of a leopard and adding a touch of elegance to its appearance.
$10+10$ A Calico lynx / bobcat with a calico pattern, featuring patches of different colors such as black, brown, and beige, creating a unique and eye-catching mosaic on its coat.
$10+10$ A Ticked lynx / bobcat exhibiting a ticked pattern, with individual hairs displaying alternating bands of light and dark colors, resulting in a subtly textured and speckled appearance.
$10+10$ A Brindle lynx / bobcat with a brindle pattern, characterized by irregular streaks and stripes of dark fur on a lighter background, adding depth and dimension to its coat.
$10+10$ A Banded lynx / bobcat displaying bands of color across its body, with distinct horizontal stripes or bands of darker fur breaking up its coat into segments.
$10+10$ A Patchwork lynx / bobcat with a patchwork pattern, featuring irregular patches of different colors and shapes distributed across its fur, creating a unique and eclectic appearance.
$10+10$ A Mackerel lynx / bobcat with a mackerel pattern, characterized by vertical stripes along its body, resembling the markings of a tiger, adding a sleek and dynamic quality to its coat.
$10+10$ A Tortoiseshell lynx / bobcat with a tortoiseshell pattern, featuring a mix of black, orange, and brown patches scattered across its fur in a unique and intricate design.
Lynx (seed 0)
Lynx (seed 1)
Bobcat (seed 0)
Bobcat (seed 1)
Striped lynx
Tabby lynx
Marbeled bobcat
Rosette bobcat
Figure 3: Examples of generated textures given various text prompts. The prompt in each row is used either with lynx or bobcat terms. The first column shows how many times each prompt (in both lynx and bobcat variant) was used with different seeds. In total, 300 textures were generated using 28 various text prompts.
@@
Environment: Our goal is to generate synthetic data that closely resembles real images. To achieve this, we use highly realistic, freely available assets from Unity’s Book Of The Dead, including terrain, trees, logs, etc. We selected several real-world scenes as references and recreated four photorealistic scenes based on them. To enhance data variability, each scene is generated in multiple versions, differing in tree types, grass, and snow coverage. See comparison of real and synthetic scenes in Figure 4.
@@ 3 Data Records
The CzechLynx dataset includes real camera trap photographs and synthetic samples of the Eurasian lynx (Lynx lynx), organized around three computer vision tasks: individual identification, pose estimation, and instance segmentation. It spans more than 15 years and covers two regions in Central Europe (Southwest Bohemia and the Western Carpathians) and contains over 40,000 manually verified and labeled images. Detailed statistics on image counts, identity coverage, and geographic distribution are listed in Table 4. All images are in JPEG format, with metadata provided in structured JSON and CSV files. The metadata includes bounding boxes, 20-point anatomical keypoints, individual identity labels, observation timestamps, and camera trap locations, to name a few. The folder structure is task-specific, with separate directories for real and synthetic data, further divided by task (e.g., pose/, reid/, segmentation/) and corresponding annotations. To allow standardized development and evaluation for all use-cases, i.e., instance segmentation, animal re-identification and pose estimation, we provide multiple sub-datasets and appropriate training / validation / test splits. All are described in detail below. The dataset is available on Kaggle.
@@
Metadata: Additional metadata are provided for a vast majority of images in the CzechLynx dataset and contain info about the origin of the data, temporal context (e.g., observation date, relative age since first sighting, and encounter sequence), spatial context (e.g., 10$\times$10 km grid-cell code, location and trap identifiers, and GPS coordinates), phenotypic annotation, and dataset partitioning flags (i.e., geo-aware, time-open, and time-closed splits). These annotations support flexible filtering and grouping by identity, time, space, and experimental split, thereby enabling reproducible, domain-aware evaluation of re-identification, pose estimation, and segmentation models. See Table 5 for a detailed description of all metadata fields.
Source # Images # Observations # Individuals Localities # Period
FoE CZ – The Western Carpathians 17,689 5,262 69 347 2009–2024
FoE CZ – Southwest Bohemia   6,442 3,374 82 74 2015–2023
Šumava National Park 13,309 9,296 106 322 2016–2024
Total 37,440 17,932 219 743 2009–2024
Table 4: Summary of data from individual sources. The FoE CZ – Southwest Bohemia and Šumava National Park datasets overlap, resulting in 38 shared individual identities.
@@
Source The provider of the data. The string beskydy, sumava, or nps correspond to FoE CZ - The Western Carpathians, FoE CZ - Southwest Bohemia, and Šumava National Park, respectively
Unique name The unique identification of Lynx Lynx individual. The format is lynx_<integer> .
Path The relative path to the file in the dataset.
@@
Relative age The age derived from the difference between the actual date and the first observation of the individual in the dataset.
@@
Coat pattern Describes the lynx’s coat pattern where available with values marbled and spotted. Present in over one-third of the images.
Cell code This column contains 10 km × 10 km grid-cell identifiers in the ETRS89-LAEA (EPSG:3035) pan-European coordinate system. Each entry has the form 10kmE<easting_index>N<northing_index>
Location The string identifier of the location.
Trap ID Unique identification of camera trap.
Geo-aware split Train / test split. The images from one location belong exclusively to either.
Split Time-open Train / test split. The individuals unseen in the train split are included in the test split.
Split Time-closed Train / test split. All individuals in the test split are included in the train split.
Latitude, Longitude WGS84 coordinates representing the center of the 10$\times$10 km grid cell where the observation occurred.
Table 5: Metadata fields and their definitions. For almost all images in the CzechLynx dataset, we include info about provenance, identity, temporal context, spatial context, phenotypic annotation, and dataset-split assignment.
@@ 3.1 Task-specific subsets
The CzechLynx dataset is organized into three subsets specifically tailored for (i) individual re-identification, (ii) animal pose estimation, and (iii) instance segmentation. Each subset is based on a different set of data; however, they largely overlap. The instance segmentation set includes the majority of data, where each image is paired with a pixel-level mask outlining the lynx’s body, enabling the training of segmentation models. For re-identification, we include only images in which individual coat patterns are clearly visible and where human experts have confirmed identities. The pose estimation subset is slightly smaller due to the labor-intensive annotation process, but also includes challenging samples with unknown identities, such as rare poses, occlusions, and low-visibility conditions. Where appropriate, the image data are accompanied by metadata, such as location, time of observation, viewpoint, and identity, to support model development and evaluation across various ecological scenarios.
Note: Due to the risk of revealing exact animal locations from background details, only the individual re-identification subset with cropped animals is publicly available. The pose estimation and instance segmentation subsets are available upon request111 Access to the dataset requires signing a Non-Disclosure Agreement (NDA) with Friends of the Earth, Czech Republic due to animal safety concerns.
Re-identification: This subset includes 37,400 manually annotated images of 219 individual Eurasian lynxes and 100k synthetic images. The real images were continuously collected for more than 15 years and originate from two primary sources: Hnutí Duha (at Beskydy and Šumava locations) and Šumava National Park (at the Šumava location). Only images with clearly visible flank, leg, or facial patterns (usually used in expert-based manual identification) were selected. The subset includes spatial and temporal metadata to allow the evaluation of identity recognition models under a large domain shift. It also supports evaluation protocols using geo-aware and time-aware splits.
Animal-pose [upon-request]: The pose estimation subset consists of around 15k real images and over 100k synthetically generated images, both annotated with up to 20 keypoints per individual. Due to the scarcity of naturally occurring diverse poses in camera-trap images, the real subset primarily contains side-view walking postures. To counterbalance this, the synthetic subset was generated using a custom Unity-based pipeline that incorporated diverse animations, viewpoints, and textured identities. This enables evaluation of generalization across both pose diversity and realism. The annotations are provided in a standardized COCO-format JSON file. Only images with a single individual are considered.
Instance Segmentation [upon-request]: This subset includes around 40k images annotated with binary instance masks delineating the full body of each lynx. Real images are manually labeled using CVAT, while synthetic images include automatically generated pixel-perfect masks derived from 3D rendering passes. The segmentation masks enable training of detection models capable of distinguishing the lynx from complex, natural backgrounds, including snow, forest undergrowth, and shadows. Each sample includes the associated mask in RLE format, aligned spatially with its corresponding image.
@@ 3.2 Predefined Splits
To support robust evaluation under real-world constraints, the CzechLynx dataset is partitioned into three distinct data splits: one geo-aware and two time-aware. Each split is designed to test model generalization across spatial or temporal domains and is provided with a clear training-validation-test split that reflects practical conservation and ecological monitoring scenarios. The summary of sample counts and individual identities per split is presented in Table 6.
Geo-aware split: This split is designed to evaluate spatial generalization. By selecting training data collected from the Hnutí Duha (Beskydy) region and testing on data from the geographically distinct Šumava National Park. The individuals in the training and test sets are disjoint, ensuring that the evaluation reflects out-of-distribution performance across separate ecological contexts. This setting simulates the challenge of deploying models trained in one region to monitor populations in another, which is highly relevant in wider scale (e.g., European) conservation monitoring.
Time-aware open-set split: In this split, data is divided based on time, with the majority of individuals in the test set overlapping with those in the training set. However, test images are drawn from a later time period, thus simulating realistic scenarios where individual appearances might have changed due to seasonal variation, aging, or camera angle. This requires models to generalize to natural visual drift and environmental changes.
Time-aware closed-set split: This split also divides data based on time but ensures that all individuals in the test set are present in the training set. This configuration focuses on evaluating how well models maintain performance over time when all identities are known during training. It reflects practical deployments in ongoing monitoring programs where the same individuals are observed across multiple seasons or years.
# images # identities # locations
training test training test training test
Geo-aware-open 19,751 17,689 150 69 396 347
Time-aware-open 27,959 9,481 197 79 667 292
Time-aware-closed 28,631 8,809 219 213 699 497
Table 6: Number of identities and images in three provided splits; two time-aware and one geo-aware.
4 Technical Validation
To ensure the technical quality of the CzechLynx dataset, we performed a series of validation steps focused on the consistency, accuracy, and usefulness of the data for both computer vision and ecological research.
@@ 4 Technical Validation
Data Consistency and Identity Validation. The re-identification subset includes only images where individual identity could be confirmed by at least two independent observers, based on distinct coat patterns on limbs and flanks. Any ambiguous cases were excluded. This process ensured a high level of reliability in the identity labels across the dataset.
Synthetic Data Comparison. To address pose diversity limitations, a synthetic subset was generated using a 3D lynx model and the Unity engine. We performed side-by-side comparisons of real and synthetic images (see Figure 4) and confirmed that the synthetic data closely mirrors real-world camera trap imagery. The synthetic samples also include complete metadata, such as keypoints, segmentation masks, and identity annotations.
Data Distribution and Coverage. The dataset spans more than 15 years and covers over 219 individual lynx across two populations and multiple regions. Images were collected from over 743 locations, with the majority sampled systematically in a 10$\times$10 km grid. This ensures representative coverage of habitats and individual variation. The dataset also includes time-aware and geographic splits, which allow robust open-set and closed-set evaluation.
@@
The dataset and baseline have been made available through the WildlifeDatasets library. We also include baseline notebooks on the Kaggle in a Code section.

### v2

@@ 2.5 Synthetic Data Generation
Synthetic data has proven valuable in improving performance in many scenarios, especially when real-world data is scarce, difficult to collect, or lacks sufficient diversity [42, 12, 43]. In the context of individual animal identification or pose estimation, however, generating synthetic data that accurately captures the complexity of real-world conditions remains a significant challenge. Despite its usefulness, existing methods for generating synthetic animal pose data often struggle with realism in appearance, motion, and environmental context, which can lead to reduced accuracy in downstream applications [44, 45, 46]. To complement the real data and broaden the variability in both poses and textures, a synthetic dataset using rendered 2D images of a 3D lynx model was created. The pipeline, built with the Unity game engine, enables the creation of highly realistic synthetic samples representing multiple individuals of the Lynx lynx species across a wide range of conditions. By leveraging detailed environmental modeling, including vegetation, terrain, and lighting, along with lifelike animations and diverse camera setups, the production of high-fidelity synthetic images that resemble real-world scenarios was ensured.
Data generation pipeline: The proposed pipeline generates synthetic animal images with rich metadata, including precise pose, segmentation masks, and individual IDs. The process starts with building a 3D environment, populated with a rigged lynx model and realistic scenery. Four control points are manually placed in each scene, some of which are outside the camera’s field of view, and the animal is animated to move between them on predefined paths. Along each path, a random set of $n$ “stop” points is sampled. At every stop, both the scene and the model are modified: environment variants (e.g., tree type, grass, snow coverage) and small changes to the camera viewpoint (e.g., rotation) are applied, while the model’s orientation and animation state (e.g., walking, sitting) are adjusted. If multiple textures corresponding to a specific identity are available, one is randomly selected for the current frame. For each valid view, a 2D snapshot is recorded together with keypoints, bounding box, and the active texture (identity), and two samples are stored per stop (before and after modifications). Frames in which most of the model lies outside the camera’s view are discarded. This procedure produces a diverse set of images that mimics real camera trap conditions. See pseudo code in Algorithm 1 for a simple overview.
Algorithm 1 Synthetic Data Generation Pipeline
1:Blank 3D mesh of animal; set of textures (each with probability $p_{t}$); set of animations (each with probability $p_{a}$)
@@
9:     With probability $p_{t}$, select a texture for the 3D mesh
10:     With probability $p_{a}$, trigger auxiliary animations
11:     Capture and store a 2D snapshot every NN-th frame
@@
Animal model: A textured model of Lynx lynx is used as the animal model. The model’s skeletal structure consists of 20 keypoints and follows the Animal Pose dataset specification , therefore, allows for training on both real and synthetic datasets using the same model and enables direct performance comparisons. Since the selected model included only a walking animation, an additional sitting animation was manually developed to increase pose diversity. With proper setup, other animal models can also be used.
Textures: To increase appearance diversity between synthetic individuals, an additional set of coat textures was generated using the Paint-It diffusion-based text-to-texture model [47]. Such a model enables the creation of textures from short text descriptions. As illustrated in Figure 3, the variability of the generated textures depended strongly on the choice of prompt rather than on the random seed. Therefore, a library of 28 prompts was defined by combining species terms (e.g., lynx, bobcat) with descriptors of common feline coat patterns (e.g., spotted, tabby, marbled) and additional details. Each prompt was sampled with multiple random seeds. Despite increasing visual variability, this approach also has some practical limitations, as the diffusion model sometimes produces artifacts or unrealistic color combinations. To mitigate this, the set of prompts was iteratively refined: prompts that consistently led to low-quality textures were replaced, whereas prompts that produced stable, realistic patterns were kept. Using the final prompt set and seed combinations, 299 textures were generated and used to represent different synthetic individuals in the simulation.
Figure 3: Examples of generated textures and given text prompts. The first two rows show textures generated from the simple prompts lynx and bobcat with different random seeds, illustrating that changes in the seed lead to only minor visual differences for a fixed prompt, while changing the prompt already alters the overall color and pattern. The third row shows an example of a more complex prompt that results in a noticeably different texture. The first column indicates how many textures were generated for each prompt (across both lynx and bobcat variants and multiple seeds). In total, 299 textures were generated using 28 different text prompts.
@@
Environment: The synthetic data were designed to resemble real camera trap images closely. To achieve this, highly realistic, freely available assets from Unity’s Book Of The Dead were used, including terrain, trees, logs, and other vegetation. Several real-world scenes were selected as references, and four photorealistic scenes were recreated based on them. To enhance data variability, each scene was rendered in multiple versions that differed in tree types, grass density, and snow coverage. A comparison between real and synthetic scenes is shown in Figure 4.
@@ 3 Data Records
The CzechLynx dataset [30] includes real camera trap photographs and synthetic samples of the Eurasian lynx (Lynx lynx), organized around three downstream tasks: (i) individual identification, (ii) pose estimation, and (iii) instance segmentation. The main part of the dataset, consisting of 39,760 manually verified and labeled camera-trap images, is fixed, whereas the synthetic part, in practice, can be scaled to any size (for simple use, a synthetic subset with a similar number of individuals and images is provided on Zenodo). The real images span more than 15 years and come from two geographically distinct regions in Central Europe: Southwest Bohemia and the Western Carpathians. As the monitoring network expanded, the yearly volume of real images increased: fewer than 300 images were collected per year before 2012, compared to more than 5,000 images per year after 2020. Observations are recorded throughout the year, with the highest capture rates between January and March (40%) and fewer images during summer months (June–August). This seasonal pattern is consistent with standard lynx monitoring designs in Central Europe, where winter camera trapping is often preferred because animals are more detectable before and during the mating season and against snow-covered backgrounds [33, 48]. Detailed statistics are provided in Table 4.
Source Images Observations Ids Sites Locs Period
FoE CZ – The Western Carpathians 17,997 9,753 95 361 39 2009–2025
FoE CZ – Southwest Bohemia   6,822 1,957 102 79 32 2015–2023
Šumava National Park Administration 14,941 7,072 169 219 27 2016–2024
Total 39,760 18,782 319 659 86 2009–2025
Table 4: Summary of camera-trap sources contributing to the CzechLynx dataset. For each source, we report the number of images, observation events, identified individuals (Ids), camera sites, spatial locations (Locs; $\times10$ km grid cells), and sampling period. The FoE CZ – Southwest Bohemia and Šumava National Park Administration sources partially overlap geographically, therefore, sharing 47 identities and 12 grid cells.
All images are stored in JPEG format (with 90% compression), with metadata provided in a structured CSV file. To simplify access to the data and support standardized development and evaluation of downstream tasks, i.e., instance segmentation, individual identification, and pose estimation, all real data are distributed in a single package, even though not all components are required for every task. The synthetic data are provided in the same repository in a separate archive. Instead of maintaining separate annotation files for each downstream task, a single shared CSV file with all annotations and necessary information is provided. It contains one row per image and, for each task, indicates whether the record is used and in which split and subset (i.e., train or test).
@@
Metadata: The CzechLynx dataset [30] includes additional metadata about the origin of the data, temporal context (e.g., observation date, relative age since first sighting, and encounter sequence), spatial context (e.g., $\times10$ km grid‐cell code, location and trap identifiers, and GPS coordinates), phenotypic annotation, and dataset partitioning flags (i.e., geo‐aware, time-aware open, and time‐aware closed splits). These annotations support flexible filtering and grouping by identity, time, space, and experimental split, thereby enabling reproducible, domain‐aware evaluation. The majority of metadata attributes are available for all images, observation dates are available for 98.9% of images, location attributes for approximately 92%, and coat-pattern annotations for about 40%. See Table 5 for a detailed description of all metadata fields.
@@
Source Data provider. The string foe_carpaths, foe_bohemia, or snpa correspond to FoE CZ - The Western Carpathians, FoE CZ - Southwest Bohemia, and Šumava National Park, respectively.
Unique name Unique identification of Lynx lynx individual. The format is lynx_<integer> .
Path Relative path to the file in the dataset.
@@
Relative age Relative age derived from the difference between the actual date and the first observation of the individual in the dataset.
@@
Coat pattern Describes lynx’s coat pattern with values marbled and spotted.
Latitude, Longitude WGS84 coordinates of the center of the $\times10$ km grid cell containing observation.
Cell code $\times10$  km grid‐cell identifier in the ETRS89-LAEA (EPSG:3035) pan-European coordinate system. Each entry has the form 10kmE<easting_index>N<northing_index>
Location Unique location identifier. The closest geopolitical region to the center of the $\times10$  km cell.
Trap ID Unique identification of camera trap. There could be multiple in each grid cell.
Geo-aware split Train / test split. Images from sources with distinct populations belong to one or the other.
Time-aware closed split Train / test split. All individuals are included in both the training and test subsets.
Time-aware open split Train / test split. Individuals unseen in the train split are included in the test split.
Pose split Train / test split. Empty if the image is not used for pose estimation.
Mask Pixel-level instance segmentation mask, stored as a COCO-style RLE.
Pose 2D pose annotation, with up to 20 visible keypoints per individual stored as a dict {<keypoint_name>: [x, y]}; empty if no pose annotation is available.
Table 5: Available metadata and their definitions.Besides annotations, the CzechLynx dataset includes standardized information on observation identity, timing, spatial context, phenotypic annotations, and dataset splits. Spatial attributes follow a hierarchical structure: (i) WGS84 coordinates of the grid-cell centroid (coordinates), (ii) a $\times10$  km ETRS89-LAEA grid cell (cell code), and (iii) the nearest administrative region to the grid-cell center (location). For the definitions of the geo-aware, time-open, and time-closed splits, see Section 3.2.
@@ 3.1 Task-specific subsets
The CzechLynx dataset [30] is organized into three task-specific subsets designed to support the development and evaluation of (i) individual identification, (ii) pose estimation, and (iii) instance segmentation. For individual identification and instance segmentation, the same set of images with clearly visible coat patterns, for which human experts confirm the identity, is provided. Each image is paired with an identity label and a pixel-level mask outlining the lynx body, which enables the training of segmentation models while providing suitable input for re-identification. The pose estimation part is a subset of the identification/segmentation images and is smaller due to the labor-intensive annotation process. Images are selected to cover a broad range of viewpoints and behaviors and also include challenging samples with rare poses, partial occlusions, and low-visibility conditions.
Individual identification: This subset contains all real images for which individual identity can be reliably determined, together with a large synthetic complement. The real images span more than 15 years and come from two primary sources: FoE CZ (at the Western Carpathians and Southwest Bohemia) and the Šumava National Park Administration (at Šumava National Park and the Protected Landscape Area in southwestern Bohemia). All images are also part of the instance-segmentation subset, paired with a pixel-level mask of the lynx body. Spatial and temporal metadata are provided to enable evaluation of identity recognition models under large domain shifts, including protocols based on geo-aware and time-aware splits (see Section 3.2.)
Instance segmentation: This subset uses the same real images as the one for individual identification. For all images, a binary instance mask delineating each lynx’s full visible body is provided. The segmentation masks enable training and evaluation of models that separate lynx from complex natural backgrounds, including snow, forest undergrowth, and shadows. All masks are provided in COCO format using run-length encoding (RLE) and are spatially aligned with their corresponding images.
Animal pose: The pose-estimation subset consists of around 5k real and a large set of synthetically generated images, each annotated with up to 20 keypoints per individual (only visible keypoints are provided). The real images primarily capture side-view walking postures, reflecting the natural bias of camera trap data. Only images with a single visible individual are included, and annotations are provided in a standardized COCO format.
@@ 3.2 Predefined Splits
To support robust evaluation under real-world constraints, the CzechLynx dataset [30] is partitioned into three distinct data splits: one geo-aware and two time-aware. Each split is designed to test model generalization across spatial or temporal domains and is provided with a clear training-test split that reflects practical conservation and ecological monitoring scenarios. The summary of image counts and individual identities per split is presented in Table 6.
Time-aware closed-set split: In this split, data are divided by time and include the same individuals in both the training and test sets. Although this setup is less ecologically realistic, since wild populations are rarely closed, it provides a controlled evaluation scenario comparable to standard benchmarks used in machine learning. The data are separated using clear time cutoffs, with no samples captured within the same encounter or in temporally neighboring periods appearing in both the training and test sets, preventing train-to-test data leakage.
Time-aware open-set split: In this split, data are also divided by time, with most individuals (82) in the test set also present in the training set. However, the test images come from later time periods, separated from the training data by season. This setup reflects realistic long-term monitoring scenarios in which populations naturally change over time (e.g., new individuals are born, some individuals die, and others may age or change appearance). The split, therefore, evaluates a model’s ability to generalize across temporal variation and population drift. The split was done using a clean time-cutoff strategy to prevent data leakage, ensuring that no images from neighboring seasons appear in both training and test sets.
Geo-aware split: This split is designed to evaluate spatial generalization. Training data come from the The Western Carpathians, while the testing comes from geographically distinct southwest Bohemia. The individuals in the training and test sets are completely disjoint, ensuring that no data leakage occurs. This setup reflects an ecological scenario where re-identification models will be deployed in a new area without locally labeled data, which tests generalization and transferability to different populations of the same species, but observed in other regions.
# of images # of identities # of sites # of locations
training test training test training test training test
Geo-aware-open 21,763 17,997 224 95 298 361 47 39
Time-aware-open 27,587 12,173 275 126 565 313 82 63
Time-aware-closed 27,836 11,924 319 319 603 464 83 77
Table 6: Number of identities, locations, sites, and images across three provided splits; two time-aware and one geo-aware. All splits were created to prevent training-to-test data leakage, i.e., samples that are close in time and/or location will never appear in both the training and test subsets. In the open-set split, 44 new identities appear in the test set.
Technical Validation
To ensure the technical quality of the CzechLynx dataset [30], we performed a series of validation steps focused on the consistency, accuracy, and usefulness of the data for both computer vision and ecological research.
@@
Data Consistency and Identity Validation. The subset for individual identification includes only images where individual identity could be confirmed by at least two independent observers, based on distinct coat patterns on limbs and flanks. Any ambiguous cases were excluded. This process ensured a high level of reliability in the identity labels across the dataset.
Synthetic Data Comparison. To address pose diversity limitations, a synthetic subset was generated using a 3D model and the Unity engine. We performed side-by-side comparisons of real and synthetic images (see Figure 4) and confirmed that the synthetic data closely mirrors real-world camera trap imagery. The synthetic samples also include similar metadata and annotations, such as keypoints and segmentation masks.
Data Distribution and Coverage. The dataset spans more than 15 years and covers 319 individual lynx across two populations and multiple regions. Images were collected from 659 sites, with the majority sampled systematically in a $\times10$ km grid. This ensures representative coverage of habitats and individual variation. The dataset also includes time- and geographically-aware splits, enabling robust open- and closed-set evaluation.
@@
Code to load and work with the CzechLynx dataset [30], including baseline training and evaluation scripts, is available in the open-source WildlifeDatasets repository. Executable baseline notebooks are provided on Kaggle in the Code section associated with the dataset. The Unity-based pipeline and diffusion-based tools used to generate the synthetic images and annotations are available in a separate repository, WildlifeSynthetic.
Data Availability
The full dataset is publicly available on Zenodo. A mirrored copy of the images and annotations, together with example notebooks, is also hosted on Kaggle.
Author contributions
Conceptualization and Methodology: Lukas Picek.
Software: Lukas Picek, Miroslav Jirik, Jakub Straka, and Vojtech Cermak.
Data Acquisition: Miroslav Kutal, Elisa Belotti, Luděk Bufka, Martin Duľa, Rostislav Dvořák, Michal Bojda,Václav Kocourek, Josefa Krausová, Jíří Labuda, Luděk Toman, Martin Váňa
Data Curation: Lukas Picek, Miroslav Jirik, and Jakub Straka.
Writing – Original Draft: Lukas Picek, Miroslav Kutal, Martin Duľa, Miroslav Jirik, Jakub Straka, Vojtech Cermak, Elisa Belotti, and Josefa Krausová.
Writing – Review & Editing: All authors read and approved the final manuscript.
Competing Interests
The author(s) declare no competing interests.